For circle H, JN = 3, NK = x, LN = 2, and NM = 6. Solve for x.



1
Expert's answer
2021-07-08T09:21:34-0400

As per the given question,

NJ = 3,

NK = x,

LN = 2,

And

NM = 6.



when two chords intersects each other, then the product of their segments is equal.

JN x NK = LN x NM

Now, substituting the values,

"\Rightarrow 3\times x=2\times 6"


"\Rightarrow x = \frac{12}{3} = 4"

"\Rightarrow x = 4"